OpenStudy (anonymous):

Are you ready for adventure? You bet you are! For this adventure, you and world renowned Professor Sherlock McMerlock are traveling to the Lost Island of Laplaya. When the boat arrives on the island shore, you and Professor McMerlock disembark on your adventure. You trudge through the jungles and arrive at three impressively large doors. About eye level on each door is an intricately carved keyhole. Directions are scratched into the wood above each keyhole.

OpenStudy (anonymous):

To open these doors, you must speak three functions in standard form. One function, f(x), with two real rational solutions. One function, g(x), with two real irrational solutions. One function, h(x), with two complex solutions. Create these three functions and explain to Professor McMerlock how you know these functions meet each condition. Remember, he is a Professor so use complete sentences. (Hint: Make sure that the b is even on g(x).)

OpenStudy (anonymous):

@phi

OpenStudy (phi):

the number of solutions a polynomial has is equal to the DEGREE (highest exponent on the variable). You want an equation with the highest exponent on x to be 2

OpenStudy (phi):

see http://www.teacherschoice.com.au/Maths_Library/Algebra/Alg_6.htm and http://www.regentsprep.org/Regents/math/algtrig/ATE3/discriminant.htm for f(x), you want the discriminant to be a perfect square (4, 9, 16, 25, etc) for g(x), you want the discriminant to not be a perfect square... lots of choices for h(x), you want the discriminant to be negative. Each function will have the form ax^2 + bx + c where you pick the 3 numbers a, b and c so that b*b - 4*a*c fits the criteria

OpenStudy (anonymous):

4x^2 + 9x + 16 would work for f(x)?

OpenStudy (anonymous):

Alright I finished my other paper so now I can focus on this one and get it done :)

OpenStudy (phi):

let's see a=4, b=9, c=16 b*b - 4*a*c 9*9 - 4 *4*16 81 - 16*16 81-256 -175 First, it is negative. That means complex (when you take the square root of a negative number, you get complex numbers) also 175 is not a perfect square 13*13 is 169 and 14*14 is 196 the square root of 175 will be between 13 and 14.

OpenStudy (anonymous):

hmm so we have to come up with a better set to get a perfect square...

OpenStudy (anonymous):

Any suggestions?

OpenStudy (phi):

one way is work backwards say you want 36 as the perfect square pick a to be a nice number, like 1! b*b - 4*a*c = 36 a is 1, so b*b -4*c = 36 b*b = 36+4*c now play around with numbers for c

OpenStudy (anonymous):

hmm.. How about 1 for c lol?

OpenStudy (phi):

not that simple, because you have b^2 = some number you need to take the square root of both sides b= sqr(number) so you want 36+4c to be a perfect square list the squares 49, 64, 81, 100 subtract 36 from each if the remainder can be divided by 4, that is a winner

OpenStudy (anonymous):

how about 64?

OpenStudy (phi):

if 4c was 64, what is 36+4*c ?

OpenStudy (anonymous):

36 + 64 = 100

OpenStudy (anonymous):

or 64 - 36 = 28 /4 = 7?

OpenStudy (anonymous):

100 - 49 = 51?

OpenStudy (anonymous):

/4 is 12.75

OpenStudy (phi):

let's start over a=1, we want b^2 - 4ac = 36 (36 is a perfect square) b^2 -4c = 36 b^2 = 36+4c we want 36+4c to be a perfect square. 36+64 = 100, which is good that means we want 4c=64 what is c ?

OpenStudy (anonymous):

c = 16

OpenStudy (phi):

b^2 = 36+4*16= 32+64= 100 b= 10 now double check: b^2 - 4 a c with a=1, b=10, c=16 10*10 - 4*1*16 100 - 64 36 and when we take the square root of 36, we get 6. so \[ f(x)= x^2 +10x +16 \] is a winner.

OpenStudy (anonymous):

lol sweet :D

OpenStudy (anonymous):

next lol

OpenStudy (anonymous):

One function, g(x), with two real irrational solutions.

OpenStudy (phi):

it just occurred to me there is an easier way: we want a quadratic with nice roots, so pick them: x= -2 and x=-8 (these are the roots we would get for h(x)) x= -2 add +2 to both sides x+2=0 similarly x+8= 0 (x+2)(x+8) = 0 multiply this out x^2 +10x +16

OpenStudy (phi):

g(x) is easy you have b^2 = 36+4c (with a=1) pick any c that does not give a perfect square for 36+4c

OpenStudy (anonymous):

uhh 3 lol?

OpenStudy (phi):

try it (I am sure it works)

OpenStudy (phi):

what is b^2 - 4 a c with a=1 , c= 3, b =10 ?

OpenStudy (anonymous):

10^2 - 4(1)(3) so... 100 - 12? which is 88

OpenStudy (anonymous):

88/4 = 22

OpenStudy (phi):

no divide by 4 the roots are \[ x= \frac{-b±\sqrt{b^2 - 4 a c}}{2a} \] you just found b^2 - 4 a c = 88 so the roots are \[ x= \frac{-10±\sqrt{88}}{2} \] you can factor 88 into 4*22 and simplify sqr(4)*sqr(22) to 2*sqr(22) \[ x= \frac{-10±2\sqrt{22}}{2} \] divide by 2 into both terms up top \[ x= -5 ± \sqrt{22} \]

OpenStudy (phi):

for the last part you want b^2 - 4 a c to be negative just make a*c bigger than b^2 i.e. pick a smaller number for b, and big numbers for a and c

OpenStudy (anonymous):

brb phone

OpenStudy (anonymous):

I'm back